# CL-Polyp: A Contrastive Learning-Enhanced Network for Accurate Polyp Segmentation

### v1

@@ Abstract
Accurate segmentation of polyps from colonoscopy images is crucial for the early diagnosis and treatment of colorectal cancer. Most existing deep learning-based polyp segmentation methods adopt an Encoder-Decoder architecture, and some utilize multi-task frameworks that incorporate auxiliary tasks such as classification to enhance segmentation performance. However, these approaches often require additional labeled data and rely on task similarity, which can limit their generalizability. To address these challenges, we propose CL-Polyp, a contrastive learning-enhanced polyp segmentation network. Our method leverages contrastive learning to improve the encoder’s ability to extract discriminative features by contrasting positive and negative sample pairs derived from polyp images. This self-supervised strategy enhances visual representation without requiring additional annotations. In addition, we introduce two lightweight and effective modules: the Modified Atrous Spatial Pyramid Pooling (MASPP) module for better multi-scale feature fusion, and the Channel Concatenate and Element Add (CA) module to fuse low-level and upsampled features for improved boundary reconstruction. Extensive experiments on five benchmark datasets—Kvasir-SEG, CVC-ClinicDB, CVC-ColonDB, CVC-300, and ETIS—demonstrate that CL-Polyp consistently outperforms state-of-the-art methods. Specifically, it improves the IoU metric by 0.011 and 0.020 on the Kvasir-SEG and CVC-ClinicDB datasets, respectively, validating its effectiveness in clinical polyp segmentation tasks.
@@ 1 Introduction
Colorectal cancer (CRC) is the third most common cancer globally, with a high incidence of colorectal adenomatous polyps being the primary manifestation[1]. In clinical practice, colonoscopy is the gold standard for identifying diseased tissues in the gastrointestinal tract. Accurate segmentation of polyps from colonoscopy images can provide essential information for treating colorectal cancer.
With the continuous development of the field of computer vision, the method of image semantic segmentation is widely used in polyp segmentation, which simplifies the tedious work of clinicians. In addition, it can better assist medical image analysis and provide more valuable information for clinical practice and pathological research.
Polyp segmentation is also a challenging medical image segmentation task, mainly due to the characteristics of colonoscopy, gastrointestinal tract, and polyps[2]. It mainly includes: 1. Irregular size and shape of polyps; 2. The color of polyps is similar to that of the surrounding intestine, so it is difficult to divide polyps from surrounding mucosa or even residual digestive matter, which also leads to the difficulty of separating the edges of polyps. 3. Polyp segmentation datasets are often small due to the high cost of manual annotation. Existing deep learning methods are gradually solving these problems.
@@
Recently, polyp segmentation methods that use CNN and/or Transformer and mainly adopt the architecture of Encoder-Decoder have achieved great success. The Encoder-Decoder architecture has been widely used in medical image segmentation since U-Net, and almost all polyp segmentation methods use this structure. To make it easier to understand, we divide such Encoder-Decoder architectures into three types according to the emphasis of the model structure (or the relative complexity of the Encoder and Decoder).
Architecture with equal emphasis on Encoder and Decoder, such as U-Net[5], U-Net++[6], and ResU-Net ++[7] with a U-shaped structure, usually under the Encoder and Decoder, uses a similar design. It transmits the features of different Encoder layers to Decoder through skip connection, generally using element addition or channel concatenation. The information obtained from the various layers of the encoder is as crucial as the information recovered from the decoder. This paper presents a constrained nonnegative matrix factorization (NMF)-based method for hyperspectral image dimensionality reduction [8].
@@
Focusing on the structure of the encoder, the design of CNN and/or Transformer is usually used to extract features, and the simple form of the decoder is used to recover detailed information.. Such as DeepLabV3+ [13], TransFuse [14], and Polyp-Pvt [15]. I think the focus of the model is the extraction of Encoder features.DeepLabV3+ uses ResNet [16] to extract features. Polyp-Pvt uses PvT[17] to extract features. TransFuse uses CNN and Transformer as Encoder to obtain better features.
@@
In order to improve the accuracy and robustness of medical image segmentation, a medical image segmentation method based on improved convolutional neural network is proposed [22]. A deep learning-based method called SR-Net combines a bidirectional recurrent convolutional network (S-Net) for inter-slice correlation and a refinement network (R-Net) for intra-slice detail enhancement, along with deformable convolutions and data consistency operations, to improve the quality and speed of undersampled multislice MR image reconstruction [23]. This paper proposes a lightweight multi-view hierarchical split network (MVHS-Net) for brain tumor segmentation in MRI images, which achieves high accuracy with significantly reduced computational complexity compared to existing methods [24]. This paper proposes a lightweight CNN-MLP hybrid model called RMMLP, which uses rolling tensors and matrix decomposition to effectively capture local and global features for accurate skin lesion segmentation, achieving state-of-the-art results with fewer parameters compared to Transformer-based methods [25]. This paper proposes a novel semi-supervised medical image segmentation method that integrates a dual-teacher architecture—combining CNN and Transformer models—with uncertainty-guided training to effectively leverage both labeled and unlabeled MRI data, achieving superior segmentation performance using limited annotations [26]. This paper proposes a lightweight 3D hippocampus segmentation method called Light3DHS, which combines a multi-scale convolutional attention module and a 3D MobileViT block to effectively fuse local and global features, achieving superior segmentation performance with fewer parameters and lower computational cost on multiple public datasets [27].
@@
As the demand for early diagnosis and treatment of colorectal cancer increases, the accuracy of polyp segmentation becomes particularly important. Currently, traditional deep learning methods have achieved some success in segmentation tasks, but they still face challenges such as small dataset sizes, high annotation costs, and the diverse shapes of polyps. These factors make it difficult for models to achieve ideal performance in practical applications. Therefore, exploring new learning methods to enhance segmentation accuracy, especially through contrastive learning to improve the model’s sensitivity to polyp features, has become a key motivation for this research. To address the above challenges, we propose adding a contrastive learning-related visual representation learning task in training the polyp segmentation model, called CL-Polyp. Contrastive learning is a self-supervised learning approach that emphasizes learning representations by contrasting positive and negative pairs. In the context of polyp segmentation, contrastive learning can enhance the model’s ability to discern subtle differences between polyp and non-polyp regions. By maximizing the similarity between positive samples—such as different views or augmentations of the same polyp—and minimizing the similarity between negative samples, the model can develop a more robust feature representation. This capability is particularly beneficial in medical imaging, where variability in polyp appearance can complicate segmentation tasks. Specifically, we add a contrastive learning task to the polyp segmentation network. According to the classification standards of the number and size of polyps in TGANet, Positive and negative samples for the contrastive learning task are selected and subjected to reasonable data augmentation. The polyp image is passed through the encoder to extract semantic information, and the encoder with momentum update is used to extract the information of positive and negative sample pairs. The whole Encoder-Decoder model is optimized by optimizing the loss function of contrastive learning and the segmentation loss function.
@@
CL-Polyp is similar to the multi-task segmentation model but different from the specific multi-task model, the model can implicitly learn features to optimize the segmentation results through contrastive learning, and better visual representation can be obtained. In addition, to better use the features in Encoder, we also modify the SOTA model DeepLabV3+. We propose the MASPP module, which can better fuse the deep multi-scale features. For recovering polyp boundary information, we presented the channel Concate and element Add (CA module) to make a better fusion between the low-level feature and upsampling feature, which can reconstruct the polyps’ boundary information. Finally, we validate our model on five challenging polyp segmentation datasets: Kvasir-SEG, CVC-ClinicDB, CVCColonDB, CVC-300, and ETIS. Specifically, on the Kvasir-SEG dataset, our method can achieve a mean Dice of 0.918 and an IoU of 0.864. On the CVC-ClinicDB dataset, our approach can perform a mean Dice of 0.955 and an IoU of 0.915.
@@
  • Extensive experiments conducted on five challenging polyp segmentation datasets demonstrate that our method outperforms existing state-of-the-art approaches in both accuracy and robustness.
Next, in Section 2, we will briefly introduce the related work in the form of Fig.1. Then, we introduce our primary method in detail in Section 3, and the experimental session is in Section 4. Finally, we make a conclusion and future work in Section 5.
@@
Figure 2: The Encoder-Decoder structure of the polyp segmentation model usually consists of three parts: Encoder, Decoder, and Skip Connection. The relative sizes of the Encoder and Decoder are expressed as structural emphasis or relative complexity. Graph a shows that Encoder and Decoder are equally important; Graph b shows that Encoder is more critical than Decoder; Graph c shows that the importance of the Decoder is higher than that of the Encoder; Graph d shows the multi-task model. The simple form of the model is to have an extra branch in the Encoder for classification tasks.
@@ 2.1 Polyp segmentation
In recent years, more segmentation methods based on deep learning have been proposed with broad applications and further development of deep learning. FCN[28] is the first architecture to implement semantic segmentation end-to-end. The proposed U-Net[5] is a U-shaped symmetric method that uses an Encoder-Decoder structure and recovers image segmentation details through skip connections. This simple and effective structure is widely used in medical image segmentation. Based on this method, many researchers have optimized on top of this benchmark method. We classify the structure of such encoder-decoders by the criteria of Encoder-Decoder relative size expressed as structural emphasis and/or relative complexity(As shown in Fig.2). The Encoder-Decoder architecture includes an Encoder to generate feature maps and obtain high-level semantic information, a Decoder to recover spatial information, and a Skip Connection to fuse low-level and high-level features to improve segmentation details(think of the structure of the skip connection as a part of the Decoder).
@@
2.The structure of the model focused on Encoder is shown in Fig.2b. This type of structure focuses on extracting features through the Encoder, but this does not mean that the Decoder is not essential; the Decoder is still a vital part. The Encoder uses the architecture of CNN and/or Transformer. The Transformer is supposed to be able to establish long-distance dependencies due to this property the Transformer. Encoders using transformers tend to obtain more valuable and robust features, but this approach increases computation due to how MSA is calculated. Some models use the Transformer structure as the Encoder to get better features, and through the fusion module, the features between the high and low layers are fused.
@@
DeeplabV3+[13], Polyp-Pvt[15], HarDNet-MSEG[31], and TransFuse[14] belong to this class of structures. The DeepLabV3+ model uses ResNet[16] as the Encoder model and uses the atrous convolution ASPP module to achieve the best effect on natural image semantic segmentation using only two upsampling. The HarDNet-MSEG model uses the HarDBlk module as the Encoder, which can get features Polyp-pvt and TransFuse differ from DeepLabV3+, and it uses the structure of a Transformer as Encoder. Polyp-Pvt uses PvT[17] as Encoder and proposes a cascaded feature fusion module in Decoder, which achieves good results. TransFuse combines CNN and Transformer as the Encoder to extract spatial information and context dependencies.
@@
3.The structure of the model focused on Decoder is shown in Fig.2c. This kind of structure usually uses the existing Backbone structure to extract features. Still, the focus of the structure is on the Decoder part. Compared with other models, the model focusing on the Decoder is the most. PraNet[10], SA-Net[11], MSNet[32], CaraNet[33] and GLFRNet[34] belong to this category. PraNet uses Res2Net[35] as Encoder, adds an inverted attention module to the Decoder, and combines deep supervision to focus on the boundary area of polyps, which has a good effect. SA-Net adds shadow attention in the Decoder to improve the effect of small polyps. In addition, it proposes color exchange data enhancement to solve a certain over-fitting problem. BoxPolyp[36] uses SA-Net as the baseline model and utilizes the bounding box utilization of a larger dataset, and solves a certain over-fitting problem by combining the bounding box with the model results. MSNet added MS Module in skip connection and Decoder, with multi-level and multi-stage cascaded subtraction operations, the complementary information from lower order to higher order among different levels can be effectively obtained. GLFRNet focuses on the process of skip connection and Decoder. The GFR module expands the receptive field and enriches the low-level semantic information. Then the LFR module is used for upsampling and feature fusion. A semi-supervised CT image segmentation method combines CNN and Transformer using entropy-constrained contrastive learning, which can improve segmentation accuracy by filtering unreliable pseudo-labels and leveraging both local and global feature information [37].
@@
Some lightweight structures also belong to Fig.2c. Such as FRCNet[38]. FRCnet proposes the ECC module and the PCF module. The simple CNN structure was used as the Encoder, ECC module was added to the Decoder to fuse the fusion between up-sampling, and PCF was used for context fusion.
@@ 2.2 Multi-task model
Multi-task learning can enhance the Encoder without making many modifications to the model and without increasing the computational load of the model. However, the disadvantage of multi-task models is that other tasks have specific results, which may need to be more effective for segmentation. If not, it may even have the opposite effect. At the same time, it requires more label information, which increases the workload of medical staff.
@@ 2.3 Contrastive learning
Contrastive learning is the ability to learn generalization and transfer from unlabeled data. The general method is to construct positive and negative samples for each data sample, and the core is to shorten the distance from the positive sample. At the same time, the distance between negative samples is increased. In this process, the selection of negative samples is the most important. MoCo[40] uses dynamic memory to store negative samples in momentum encoder storage. SimCLR[41] shows that large batch sizes as well as long training times are also important for contrastive learning. By introducing a stopping gradient, SimSiam[42] indicates that meaningful representations can be learned from simple network architectures without negative pairs, large batch sizes, or momentum encoders. Few researchers have used contrastive learning in polyp segmentation models, although this method is very effective.

### v2

# CL-Polyp: Contrastive Learning for Accurate Polyp Segmentation

@@ Abstract
Accurate segmentation of polyps from colonoscopy images is crucial for the early diagnosis and treatment of colorectal cancer. Most existing deep learning-based polyp segmentation methods adopt an Encoder-Decoder architecture, and some utilize multi-task frameworks that incorporate auxiliary tasks like classification to improve segmentation. However, these methods often need more labeled data and depend on task similarity, potentially limiting generalizability. To address these challenges, we propose CL-Polyp, a contrastive learning-enhanced polyp segmentation network. Our method uses contrastive learning to enhance the encoder’s extraction of discriminative features by contrasting positive and negative sample pairs from polyp images. This self-supervised strategy improves visual representation without needing additional annotations. We also introduce two efficient, lightweight modules: the Modified Atrous Spatial Pyramid Pooling (MASPP) module for improved multi-scale feature fusion, and the Channel Concatenate and Element Add (CA) module to merge low-level and upsampled features for enhanced boundary reconstruction. Extensive experiments on five benchmark datasets—Kvasir-SEG, CVC-ClinicDB, CVC-ColonDB, CVC-300, and ETIS—show that CL-Polyp consistently surpasses state-of-the-art methods. Specifically, it enhances the IoU metric by 0.011 and 0.020 on the Kvasir-SEG and CVC-ClinicDB datasets, respectively, demonstrating its effectiveness in clinical polyp segmentation.
@@ 1 Introduction
Colorectal cancer (CRC) is the third most common cancer globally, with a high incidence of colorectal adenomatous polyps being the primary manifestation[1]. In clinical practice, colonoscopy is the gold standard for identifying diseased tissues in the gastrointestinal tract. Accurate polyp segmentation in colonoscopy images provides essential information for colorectal cancer treatment.
With advances in computer vision, image semantic segmentation is now widely used for polyp segmentation, simplifying clinicians’ work. It also enhances medical image analysis and provides more valuable information for clinical practice and pathological research.
Polyp segmentation is also a challenging medical image segmentation task, mainly due to the characteristics of colonoscopy, gastrointestinal tract, and polyps[2]. It mainly includes: 1. Irregular size and shape of polyps; 2. Polyps often have similar color to the surrounding intestine, making it difficult to distinguish them from mucosa or residual digestive matter, complicating edge separation. 3. Polyp segmentation datasets are often small due to the high cost of manual annotation. Existing deep learning methods are gradually solving these problems.
@@
Recent polyp segmentation methods using CNN and/or Transformer, and mainly adopting the Encoder-Decoder architecture, have achieved significant success. The Encoder-Decoder architecture has been widely used in medical image segmentation since U-Net, and almost all polyp segmentation methods use this structure. For clarity, we categorize these Encoder-Decoder architectures into three types based on the emphasis of the model structure (or the relative complexity of the Encoder and Decoder).
Architecture with equal emphasis on Encoder and Decoder, such as U-Net[5], U-Net++[6], and ResU-Net ++[7] with a U-shaped structure, usually under the Encoder and Decoder, uses a similar design. It transmits the features of different Encoder layers to Decoder through skip connection, generally using element addition or channel concatenation. The information obtained from the various layers of the encoder is as crucial as the information recovered from the decoder. This paper presents a constrained nonnegative matrix factorization (NMF)-based method for hyperspectral image dimensionality reduction [8, 9]. Hybrid Parallel Genetic Algorithm addresses the limitations of traditional segmentation methods—such as dependence on initial values, susceptibility to local optima, and slow computation—through parallel evolution and global optimization, making it particularly suitable for automated segmentation tasks involving high-dimensional, high-resolution, and multi-modal medical images [10].
@@
Focusing on the structure of the encoder, the design of CNN and/or Transformer is usually used to extract features, and the simple form of the decoder is used to recover detailed information. Such as DeepLabV3+ [15], TransFuse [16], and Polyp-Pvt [17]. I think the focus of the model is the extraction of Encoder features. DeepLabV3+ uses ResNet [18] to extract features. Polyp-Pvt uses PvT[19] to extract features. CNN-ViT hybrid architectures have been widely adopted for image analysis[20, 21, 22, 23, 24, 25].
@@
In order to improve the accuracy and robustness of medical image segmentation, a medical image segmentation method based on improved convolutional neural network is proposed [30, 31, 32]. A deep learning-based method called SR-Net combines a bidirectional recurrent convolutional network (S-Net) for inter-slice correlation and a refinement network (R-Net) for intra-slice detail enhancement, along with deformable convolutions and data consistency operations, to improve the quality and speed of undersampled multislice MR image reconstruction [33]. This paper proposes a lightweight multi-view hierarchical split network (MVHS-Net) for brain tumor segmentation in MRI images, which achieves high accuracy with significantly reduced computational complexity compared to existing methods [34]. This paper proposes a lightweight CNN-MLP hybrid model called RMMLP, which uses rolling tensors and matrix decomposition to effectively capture local and global features for accurate skin lesion segmentation, achieving state-of-the-art results with fewer parameters compared to Transformer-based methods [35]. This paper proposes a novel semi-supervised medical image segmentation method that integrates a dual-teacher architecture—combining CNN and Transformer models—with uncertainty-guided training to effectively leverage both labeled and unlabeled MRI data, achieving superior segmentation performance using limited annotations [36]. This paper proposes a lightweight 3D hippocampus segmentation method called Light3DHS, which combines a multi-scale convolutional attention module and a 3D MobileViT block to effectively fuse local and global features, achieving superior segmentation performance with fewer parameters and lower computational cost on multiple public datasets [37].
@@
As the demand for early diagnosis and treatment of colorectal cancer increases, the accuracy of polyp segmentation becomes particularly important. Currently, traditional deep learning methods have achieved some success in segmentation tasks, but they still face challenges such as small dataset sizes, high annotation costs, and the diverse shapes of polyps. These factors make it difficult for models to achieve ideal performance in practical applications. Therefore, exploring new learning methods to enhance segmentation accuracy, especially through contrastive learning to improve the model’s sensitivity to polyp features, has become a key motivation for this research. To address the above challenges, we propose adding a contrastive learning-related visual representation learning task in training the polyp segmentation model, called CL-Polyp. Contrastive learning is a self-supervised learning approach that emphasizes learning representations by contrasting positive and negative pairs. In the context of polyp segmentation, contrastive learning can enhance the model’s ability to discern subtle differences between polyp and non-polyp regions. By maximizing the similarity between positive samples—such as different views or augmentations of the same polyp—and minimizing the similarity between negative samples, the model can develop a more robust feature representation. This capability is particularly beneficial in medical imaging, where variability in polyp appearance can complicate segmentation tasks. Specifically, we add a contrastive learning task to the polyp segmentation network. According to the classification standards of the number and size of polyps in TGANet, Positive and negative samples for the contrastive learning task are selected and subjected to reasonable data augmentation. The polyp image is passed through the encoder to extract semantic information, and the encoder with momentum update is used to extract the information of positive and negative sample pairs. The entire Encoder-Decoder model is optimized by minimizing both the contrastive learning and segmentation loss functions.
@@
CL-Polyp is similar to the multi-task segmentation model but different from the specific multi-task model, the model can implicitly learn features to optimize the segmentation results through contrastive learning, and better visual representation can be obtained. In addition, to better use the features in Encoder, we also modify the SOTA model DeepLabV3+. We propose the MASPP module, which can better fuse the deep multi-scale features. For recovering polyp boundary information, we presented the channel Concate and element Add (CA module) to make a better fusion between the low-level feature and upsampling feature, which can reconstruct the polyps’ boundary information. Finally, we validate our model on five challenging polyp segmentation datasets: Kvasir-SEG, CVC-ClinicDB, CVCColonDB, CVC-300, and ETIS. Specifically, on the Kvasir-SEG dataset, our method achieves a mean Dice of 0.918 and an IoU of 0.864. On the CVC-ClinicDB dataset, our approach achieves a mean Dice of 0.955 and an IoU of 0.915.
@@
  • Extensive experiments on five challenging polyp segmentation datasets demonstrate that our method surpasses existing state-of-the-art approaches in accuracy and robustness.
Next, in Section 2, we will briefly introduce the related work in the form of Fig.1. Then, we introduce our primary method in detail in Section 3, and the experimental session is in Section 4. Finally, Section 5 concludes and discusses future work.
@@
Figure 2: Encoder-decoder structure variations for polyp segmentation models: (a) Balanced encoder and decoder; (b) Encoder with higher relative complexity; (c) Decoder with higher relative complexity; (d) Multi-task model with an auxiliary classification branch in the encoder.
@@ 2.1 Polyp segmentation
Recently, numerous deep learning-based segmentation methods have been proposed, leading to broad applications and further development. FCN[38] is the first architecture to implement semantic segmentation end-to-end. The proposed U-Net[5] is a U-shaped symmetric method that uses an Encoder-Decoder structure and recovers image segmentation details through skip connections. This simple and effective structure is widely used in medical image segmentation. Based on this method, many researchers have optimized this benchmark. We classify the structure of such encoder-decoders by the criteria of Encoder-Decoder relative size expressed as structural emphasis and/or relative complexity(As shown in Fig.2). The Encoder-Decoder architecture comprises an Encoder for generating feature maps and extracting high-level semantic information, a Decoder for recovering spatial information, and Skip Connections for fusing low-level and high-level features to enhance segmentation details (consider the skip connections as part of the Decoder).
@@
2.The structure of the model focused on Encoder is shown in Fig.2b. This type of structure focuses on extracting features through the Encoder, but this does not mean that the Decoder is not essential; the Decoder is still a vital part. The Encoder uses the architecture of CNN and/or Transformer. The Transformer is supposed to be able to establish long-distance dependencies due to this property the Transformer. Encoders using transformers tend to obtain more valuable and robust features, but this approach increases computation due to how MSA is calculated. Some models use the Transformer as the Encoder to obtain enhanced features, fusing high- and low-level features through a fusion module.
@@
DeeplabV3+[15], Polyp-Pvt[17], HarDNet-MSEG[41], and TransFuse[16] belong to this class of structures. The DeepLabV3+ model uses ResNet[18] as the Encoder model and uses the atrous convolution ASPP module to achieve the best effect on natural image semantic segmentation using only two upsampling. The HarDNet-MSEG model employs the HarDBlk module as the Encoder to extract features. Polyp-PVT and TransFuse differ from DeepLabV3+ by using a Transformer structure as the Encoder. Polyp-Pvt uses PvT[19] as Encoder and proposes a cascaded feature fusion module in Decoder, which achieves good results. TransFuse combines CNN and Transformer as the Encoder to extract spatial information and context dependencies.
@@
3.The structure of the model focused on Decoder is shown in Fig.2c. This kind of structure usually uses the existing Backbone structure to extract features. Still, the focus of the structure is on the Decoder part. Compared with other models, the model focusing on the Decoder is the most. PraNet[12], SA-Net[13], MSNet[42], CaraNet[43] and GLFRNet[44] belong to this category. PraNet uses Res2Net[45] as Encoder, adds an inverted attention module to the Decoder, and combines deep supervision to focus on the boundary area of polyps, which has a good effect. SA-Net incorporates shadow attention in the Decoder to improve small polyp segmentation and proposes color exchange data augmentation to address overfitting. BoxPolyp[46] uses SA-Net as the baseline model and utilizes the bounding box utilization of a larger dataset, and solves a certain over-fitting problem by combining the bounding box with the model results. MSNet adds an MS Module in the skip connection and Decoder, using multi-level and multi-stage cascaded subtraction operations to effectively obtain complementary information from lower to higher orders across different levels. GLFRNet focuses on skip connections and the Decoder. The GFR module expands the receptive field and enriches low-level semantic information. The LFR module then performs upsampling and feature fusion. A semi-supervised CT image segmentation method combines CNN and Transformer using entropy-constrained contrastive learning, which can improve segmentation accuracy by filtering unreliable pseudo-labels and leveraging both local and global feature information [47].
@@
Some lightweight structures also belong to Fig.2c. Such as FRCNet[48]. FRCnet introduces the ECC and PCF modules. Using a simple CNN structure as the Encoder, the ECC module is added to the Decoder to enhance fusion during up-sampling, while the PCF module is used for context fusion.
@@ 2.2 Multi-task model
Multi-task learning can enhance the Encoder without making many modifications to the model and without increasing the computational load of the model. However, the disadvantage of multi-task models is that other tasks have specific results, which may need to be more effective for segmentation. If not, it may even have the opposite effect. Furthermore, it requires additional label information, increasing the workload of medical staff.
@@ 2.3 Contrastive learning
Contrastive learning is the ability to learn generalization and transfer from unlabeled data. The general approach involves constructing positive and negative samples for each data point, aiming to reduce the distance to the positive sample. At the same time, the distance between negative samples is increased. In this process, the selection of negative samples is the most important. MoCo[50] uses dynamic memory to store negative samples in momentum encoder storage. SimCLR[51] shows that large batch sizes as well as long training times are also important for contrastive learning. By introducing a stopping gradient, SimSiam[52] indicates that meaningful representations can be learned from simple network architectures without negative pairs, large batch sizes, or momentum encoders. Few researchers have used contrastive learning in polyp segmentation models, although this method is very effective.